Silane, , reacts with to give silicon dioxide and water:
SiH4(g)+2 O2(g)-->SiO2(s)+2H2O(l)
A 4.51-L sample of gas at 350. mm Hg pressure and 25 °C is allowed to react with gas. What volume of gas, in liters, is required for complete reaction if the oxygen has a pressure of 428 mm Hg at 25 °C?

Volume =??? L
1
Expert's answer
2020-05-01T14:30:51-0400

pV=nRT;

n(SiH4)=pV/RT;

T=273+25=298 K

p=350 mmHg=46.66 kPa;

n(SiH4)=46.66*4.51/(8.314*298)=0.085 mol

n(O2) required is 0.085*2=0.17 mol

V=nRT/p=0.17*8.314*298/57.062=7.4 L

Volume=7.4 L